1 Add or 2 15 4 o Subtract given matrices 3-51 1-281

adding is what i need to do
Answer #1

Add or substract given matrices Let A= 14. A+B= (2+3 4+(-5 A+B - - 5+(-2) +8 A-B = 1 2 3 4 A-B -8. 1-5-1-2) o -8 CScanned wit